OpenStudy (anonymous):

what is the radical form of (6x)^3/2 ?

OpenStudy (anonymous):

\[\sqrt[2]{(6x)^3}\]

OpenStudy (anonymous):

that's not an answer choice on my paper, is that answer, without the 2, the same thing?

OpenStudy (amistre64):

or... written another way: sqrt(216x^3)

OpenStudy (anonymous):

yes, with out 2, same thing

OpenStudy (anonymous):

\[\sqrt{x} = \sqrt[2]{x}\] for example